Calculating the Current Ratio and the Quick (or Acid-Test) Ratio LoLo Lemon Company has current assets equal to $500,000. Of these, $300,000 is cash, $75,000 is accounts receivable, $125,000 is inventory, and the remainder is marketable securities. Current liabilities total $425,000. Required: Note: Round answers to two decimal places.

1. Calculate the current ratio.

2. Calculate the quick ratio (acid-test ratio).
Answer #1

1.

Current ratio = Current assets/Current liabilities

= 500,000/425,000

= 1.18

2.

Acid test ratio = (Cash + Accounts receivable + marketable securities)/Current liabilities

= (300,000 + 75,000 + 0)/425,000

= 375,000/425,000

= 0.88

kindly give a positive rating if you are satisfied with the solution. Please ask if you have any query related to the question, Thanks.